# Paper Copilot: Tracking the Evolution of Peer Review in AI Conferences

### v1

@@ 1 Introduction
The rapid growth of submissions to top-tier Artificial Intelligence (AI) and Machine Learning (ML) conferences—now exceeding 10,000 per venue annually (Yang, 2025)—has placed unprecedented pressure on peer review. To address issues of scale and transparency, many conferences have adopted open or semi-open platforms like OpenReview. Yet practices remain inconsistent: some venues publish reviews and scores, while others keep them private. This inconsistency has sparked ongoing debate around fairness, accountability, and the effectiveness of different review models  (Cortes & Lawrence, 2021).
@@
From the standpoint of conference organizers, internal dashboards provide visibility into score trends and acceptance distributions (Cortes & Lawrence, 2021). But these views are rarely made accessible to authors. During the critical post-review phase—often only one or two weeks long—authors are expected to analyze reviewer feedback, prepare rebuttals, and potentially decide whether to withdraw or revise their submissions. With scattered resources, authors must resort to scraping statistics (Sun, 2020) or aggregating posts from scattered social media. This inefficiency hinders timely decision-making and may even lead to missed opportunities for meaningful rebuttal or clarification. A system that actively collects review statistics, visualizes score dynamics, and contextualizes submissions relative to their peers would therefore offer immediate utility to thousands of authors.
@@ 2.1 Peer Review Datasets and Analysis
Bharti et al. (2022) leveraged an NLP approach to estimate reviewer confidence from review text, illustrating analysis that can be done when reviews are available. Peer Review Analysis (Ghosal et al., 2022) provided 1,199 review reports labeled at the sentence level for aspects like critique, suggestion, and sentiment, serving as a benchmark for automated review assessment. Ribeiro et al. (2021) used a private dataset of two non-disclosed conferences to perform RSP, PDP, and sentiment analysis to extract review polarities. In general, these datasets have been invaluable in enabling studies on review quality, bias, and consistency. However, each is often limited to a particular venue or a snapshot in time. We advance beyond prior datasets by unifying multiple venues over a long temporal window, thereby supporting cross-conference and longitudinal analyses (e.g., how review score distributions shift over years, or how an author’s successive submissions fare over time). It also introduces the notion of tracking review process “dynamics,” since our data include timestamps and sequences (allowing one to reconstruct the timeline of reviews and discussions for open-review conferences). This complements earlier one-off studies that examined review consistency and randomness in single years (Beygelzimer et al., 2021; Cortes & Lawrence, 2021), by providing a broader data foundation to study such effects in the aggregate.
@@ 2.3 Author Profiling and Metascience
Studying author career trajectories and publication patterns has been a long-standing interest in bibliometrics (Fan et al., 2024; De Frutos Belizon et al., 2025). Large-scale academic databases and knowledge graphs – such as Google Scholar, Microsoft Academic Graph, Semantic Scholar’s Open Research Corpus, and AMiner – have enabled analyses of collaboration networks, citation trajectories, and researcher mobility (Scholar, ). For example, AMiner (Tang et al., 2008), an academic social network mining system that compiled tens of thousands of author profiles and their affiliations over time, allowing for queries on career moves and productivity. A recent dataset from AMiner contains over 57,000 scholars’ career paths (with 42,230 affiliations) for studying academic job transitions. These tools, however, focus on publication and citation data; they generally do not incorporate the peer review stage. We brings a new dimension by linking authors to the review process outcomes of their submissions. This creates opportunities for “talent trajectory” studies in AI / ML – for instance, one could track how an early-career researcher improves their paper quality (as reflected in review scores) over successive submissions, or examine whether consistent high reviewers’ ratings correlate with future citation impact (Vinkenburg et al., 2020). Our dataset, combined with external bibliometric data, could also help identify patterns such as mentorship lineages (e.g., how students of certain groups perform in reviews) or geographic trends in research feedback (Guevara et al., 2016). In the broader context of metascience, our work contributes an open resource to empirically study questions about peer review and researcher development. This aligns with recent calls for transparency and data-driven policy in peer review (Bianchi & Squazzoni, 2022). By making our data easily accessible, we hope to facilitate further meta-research on how review processes influence and reflect the growth of the AI / ML research community.
@@ 8 Acknowledgment
I, Jing Yang, sincerely thank the global AI/ML community for their trust, contributions, and engagement with Paper Copilot. Independently built from the ground up without institutional backing, this project was developed and maintained during my spare time and fully self-funded. I’m especially grateful to the researchers who voluntarily shared data in support of transparency and collective progress.
I also wish to thank my advisor, Prof. Yajie Zhao, for fostering academic freedom. While this work was conducted independently of my advisor and lab, her support provided the flexibility to pursue it. More details can be found at : https://papercopilot.com/acknowledgment/
Finally, I gratefully acknowledge the valuable discussions, feedback, and support from Hao Sun and Samuel Holt (University of Cambridge), Markus Wulfmeier (Google DeepMind), and Prof. Nihar B. Shah (CMU), whose thoughtful input helped shape the direction of this work.
@@ Appendix C Author–Affiliation Metadata Acquisition
Venue
Years
#Papers
Author–
Affil.
Affil.–
Country
Multi-Affil.
Prompt
Tokens
Completion
Tokens
Total
Tokens

### v2

@@ 1 Introduction
The rapid growth of submissions to top-tier Artificial Intelligence (AI) and Machine Learning (ML) conferences—now exceeding 10,000 per venue annually (Yang, 2025)—has placed unprecedented pressure on peer review. To address issues of scale and transparency, many conferences have adopted open or semi-open platforms like OpenReview. Yet practices remain inconsistent: some venues publish reviews and scores, while others keep them private. This inconsistency has sparked ongoing debate around fairness, accountability, and the effectiveness of different review models  (Cortes and Lawrence, 2021).
@@
From the standpoint of conference organizers, internal dashboards provide visibility into score trends and acceptance distributions (Cortes and Lawrence, 2021). But these views are rarely made accessible to authors. During the critical post-review phase—often only one or two weeks long—authors are expected to analyze reviewer feedback, prepare rebuttals, and potentially decide whether to withdraw or revise their submissions. With scattered resources, authors must resort to scraping statistics (Sun, 2020) or aggregating posts from scattered social media. This inefficiency hinders timely decision-making and may even lead to missed opportunities for meaningful rebuttal or clarification. A system that actively collects review statistics, visualizes score dynamics, and contextualizes submissions relative to their peers would therefore offer immediate utility to thousands of authors.
@@ 2.1 Peer Review Datasets and Analysis
Bharti et al. (2022) leveraged an NLP approach to estimate reviewer confidence from review text, illustrating analysis that can be done when reviews are available. Peer Review Analysis (Ghosal et al., 2022) provided 1,199 review reports labeled at the sentence level for aspects like critique, suggestion, and sentiment, serving as a benchmark for automated review assessment. Ribeiro et al. (2021) used a private dataset of two non-disclosed conferences to perform RSP, PDP, and sentiment analysis to extract review polarities. In general, these datasets have been invaluable in enabling studies on review quality, bias, and consistency. However, each is often limited to a particular venue or a snapshot in time. We advance beyond prior datasets by unifying multiple venues over a long temporal window, thereby supporting cross-conference and longitudinal analyses (e.g., how review score distributions shift over years, or how an author’s successive submissions fare over time). It also introduces the notion of tracking review process “dynamics,” since our data include timestamps and sequences (allowing one to reconstruct the timeline of reviews and discussions for open-review conferences). This complements earlier one-off studies that examined review consistency and randomness in single years (Beygelzimer et al., 2021; Cortes and Lawrence, 2021), by providing a broader data foundation to study such effects in the aggregate.
@@ 2.3 Author Profiling and Metascience
Studying author career trajectories and publication patterns has been a long-standing interest in bibliometrics (Fan et al., 2024; De Frutos Belizon et al., 2025). Large-scale academic databases and knowledge graphs – such as Google Scholar, Microsoft Academic Graph, Semantic Scholar’s Open Research Corpus, and AMiner – have enabled analyses of collaboration networks, citation trajectories, and researcher mobility (Scholar, ). For example, AMiner (Tang et al., 2008), an academic social network mining system that compiled tens of thousands of author profiles and their affiliations over time, allowing for queries on career moves and productivity. A recent dataset from AMiner contains over 57,000 scholars’ career paths (with 42,230 affiliations) for studying academic job transitions. These tools, however, focus on publication and citation data; they generally do not incorporate the peer review stage. We brings a new dimension by linking authors to the review process outcomes of their submissions. This creates opportunities for “talent trajectory” studies in AI / ML – for instance, one could track how an early-career researcher improves their paper quality (as reflected in review scores) over successive submissions, or examine whether consistent high reviewers’ ratings correlate with future citation impact (Vinkenburg et al., 2020). Our dataset, combined with external bibliometric data, could also help identify patterns such as mentorship lineages (e.g., how students of certain groups perform in reviews) or geographic trends in research feedback (Guevara et al., 2016). In the broader context of metascience, our work contributes an open resource to empirically study questions about peer review and researcher development. This aligns with recent calls for transparency and data-driven policy in peer review (Bianchi and Squazzoni, 2022). By making our data easily accessible, we hope to facilitate further meta-research on how review processes influence and reflect the growth of the AI / ML research community.
@@ 8 Acknowledgment
We sincerely thank the global AI/ML community for their trust, contributions, and engagement with Paper Copilot. We are especially grateful to the researchers who voluntarily shared data in support of transparency and collective progress.
Specifically, the first author, Jing Yang, initiated and has been maintaining Paper Copilot largely during spare time, with primary support coming from personal effort and resources. The first author also thanks the advisor, Prof. Yajie Zhao, for fostering academic freedom and providing a supportive environment to pursue this direction. More details are available at https://papercopilot.com/acknowledgment/.
Finally, we gratefully acknowledge valuable discussions, feedback, and support from Hao Sun and Samuel Holt (University of Cambridge), Markus Wulfmeier (Google DeepMind), and Prof. Nihar B. Shah (CMU), whose thoughtful input helped shape the direction of this work.
@@ Appendix C Author–Affiliation Metadata Acquisition
Venue Years #Papers Author– Affil. Affil.– Country % Multi-Affil. Prompt Tokens Completion Tokens Total Tokens